OpenStudy (anonymous):

Rewrite the equation m1v1+m2v2=m1v1'+m2v2 a. in the case of m1=m2=m (i.e., replace m1 and m2 with just m) b. in the case of v1=v2=v c. in the case of v1'=v2'=v'

OpenStudy (anonymous):

I have no idea what this question is asking. Try to use the equation editor below to see if you can re-post the question more clearly.

OpenStudy (anonymous):

\[m _{1}\times v_{1} + m _{2} \times v _{2} = m _{1} \times v _{1}\prime + m _{2}\times v _{2}\prime \]

OpenStudy (anonymous):

Oh, okay. In that case, a should be really easy. Use the distributive property after you've made both m1 and m2 equal to m.

OpenStudy (anonymous):

How do I make m1 and m2 equal to m? Do I just plug it in?

OpenStudy (anonymous):

Yup. For example, if I have 5x, and x = 3, I have 5*3. If x = z, it's the same idea, but now I have 5*z.

OpenStudy (anonymous):

So for a. would it just look like mv1+mv2=mv1'+mv2?

OpenStudy (anonymous):

Almost. Now that you have m in both variables on both sides, can you combine them?

OpenStudy (anonymous):

But if I did that wouldn't it turn out to be 0=0 because I would have to subtract mv1 from both sides and then mv2 from both sides which would equal 0.

OpenStudy (anonymous):

Shouldn't. Because you have v1' on one side. Go ahead and show me how you would write it out. Even if you think it would end up as 0. I can point out where you're getting caught.

OpenStudy (anonymous):

mv1+mv2=mv1'+mv2 -mv1 mv2=0+mv2 -mv2 0=0

OpenStudy (anonymous):

Okay... so I think I see it. mv1' is different from mv1, so you can't subtract them.

OpenStudy (anonymous):

Have you covered calculus, or derivatives at all in your course?

OpenStudy (anonymous):

No, I'm in Algebra 2. This is in my Physics class.

OpenStudy (anonymous):

Gotcha. Okay, well v1' and v1 are two different things. You may as well treat them as completely different variables; they are not interchangeable. If they were, you'd be totally correct and everything would cancel out.

OpenStudy (anonymous):

So how do I simplify it?

OpenStudy (anonymous):

Well, now that you know that you should treat them as different variables, and that you can use the distributive property, apply those two and you'll have an answer for a. Which works like this: 5(x + 5) = 5x + 5*5

OpenStudy (anonymous):

*that was an example of the distributive property. In this case, you'll probably want to use it in reverse...

OpenStudy (anonymous):

m(v1+v2)=m(v1'+v2')?

OpenStudy (anonymous):

Perfect. Now you can cancel out the m's.

OpenStudy (anonymous):

So it's just v1+v2=v1'+v2'

OpenStudy (anonymous):

That's as far as I would take it, unless you know the value of any of those variables.

OpenStudy (anonymous):

It may not seem like a lot, but being able to eliminate a variable will save you loads of time. You want to get good and being able to recognize when you can do that. :)

OpenStudy (anonymous):

*at being able to recognize

OpenStudy (anonymous):

So, just to make sure I really know this, b. v(m1+m2)=m1v1'+m2v2' c. m1v1+m2v2=v(m1+m2) Is that all right?

OpenStudy (anonymous):

Looks like it. Double check what you "pulled out" in c, though. It shouldn't be just "v".

OpenStudy (anonymous):

c. m1v1+m2v2=v'(m1+m2)?

OpenStudy (anonymous):

There you go. :)

OpenStudy (anonymous):

I would double check with at least one other person, though - always good to have more than one perspective.

OpenStudy (anonymous):

Thanks a lot for all the help!

OpenStudy (anonymous):

You're welcome!